# Verification and Validation for Trustworthy Scientific Machine Learning

### v1

@@ 6.7 Explain and Interpret
The limited interpretability and transparency of ML models are significant challenges, especially in critical scientific applications such as medicine and aerospace engineering. The inner workings of SciML models are hard for humans to understand, possibly deterring us from trusting their predictions. This is especially problematic in settings where model-assisted decisions can significantly impact people’s lives. To address these problems, researchers must continue to develop new techniques for making SciML models more interpretable and transparent.
While many papers compare the interpretability of different methods, no universal definition of interpretability exists because interpretability is subjective and varies across scientific disciplines [60]. Linear models are often considered interpretable because coefficients show each feature’s importance in predictions. However, this interpretability breaks down with thousands of parameters [68]. Claims of superior interpretability must be thoroughly justified. This creates an opportunity for research communities to define domain-specific properties that make models interpretable, complemented by explainability methods that reveal how models make their predictions.

### v2

@@ 6.7 Explain and Interpret
The limited interpretability and transparency of machine learning (ML) models pose significant challenges in critical fields like medicine and aerospace engineering. The complexity of Scientific Machine Learning (SciML) models can hinder trust in their predictions, particularly when decisions impact lives. Unfortunately, a universal definition of interpretability remains elusive due to its subjective nature and variation across disciplines [48]. Linear models are often deemed interpretable because their coefficients indicate feature importance; however, this clarity diminishes with thousands of parameters [56]. Thus, claims of superior interpretability require thorough justification, presenting an opportunity for research communities to establish domain-specific criteria for interpretability, alongside explainability methods that clarify model predictions.